OpenStudy (anonymous):

antiderivative of 1+cos^2(x) secx dx?

hartnn (hartnn):

whats cos x sec x = .. ?

OpenStudy (anonymous):

that's my problem, I don't know how to rewrite that

OpenStudy (anonymous):

isn't it simply 1?

hartnn (hartnn):

its \(\int 1+ \cos^2 x \sec x dx\) right ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

yes so cos^2 x sec x is just cos x

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

\( \int 1+\cos x dx\)

OpenStudy (anonymous):

using def of secx right?

OpenStudy (anonymous):

so antiderivative is sinx+x

hartnn (hartnn):

+C yes

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^